# CATArena: Evaluation of LLM Agents Through Iterative Tournament Competitions

### v1

@@ 4.1 Experimental Setups
In our experiment, we employ three types of agents: (1) Minimal Agents (LLM + ADK Framework) [^2] https://github.com/AGI-Eval-Official/Minimal-CodeAgent: A baseline agent developed with the Agent Development Kit (ADK) Python toolkit. We provide essential tools, including file manipulation, bash scripting, and Python execution, for the ADK code agent. On this foundation, we integrate state-of-the-art LLMs to systematically compare their core competencies as code agents for strategy implementation. (2) Commercial Code Agents: State-of-the-art and commercial CLI-based agents (e.g., Claude Code, CodeX, Gemini-CLI, Qwen-Coder) are included for benchmarking. These agents feature advanced integration with various command-line interfaces, tools, and LLMs, resulting in enhanced overall capabilities. These agents serve as leading solutions in code agent development and provide valuable reference points for future research. (3) LLM-Player: In this control setting, LLMs directly output game moves without generating code. For each turn, the LLM receives the game rules, current state, and history, and returns the next action. This approach is specifically designed to assess the inherent strategic and reasoning capabilities of LLMs. Detailed agent parameter settings and model selections are presented in Appendix C.
Table 2: Agent Specifications and Open-Source Status.
Agent Type Agent Framework Model Agent OSS LLM OSS
Qwen3-Coder-480B (Team, 2025c)
Doubao-Seed-1.6 (Team, 2025d)
GPT-5 (OpenAI, 2025b)
Claude-4-Sonnet (Anthropic, 2025a)
Gemini-2.5-pro (Team, 2025a)
Commercial Gemini-CLI (Google, 2025) Gemini-2.5-pro (Team, 2025a)
Claude-Code (Anthropic, 2025b) Claude-4/3.7 Hybrid (Anthropic, 2025a)
CodeX (OpenAI, 2025a) GPT-5 (OpenAI, 2025b)
Qwen-Coder (Team, 2025b) Qwen3-Coder-480B (Team, 2025c)
Other LLM-Player Agents’s Corresponding LLM N/A N/A
Tournaments.
All experiments are conducted under two main tournament settings: (1) a comparison among minimal agents equipped with different LLMs ($T_{1}=6$), and (2) a comparison between the best-performing minimal agent and a set of commercial code agents ($T_{2}=5$). To reduce the impact of randomness on strategy generation, each tournament is repeated for four times, and all reported metrics are averaged over the four runs. Each tournament consists of $N=4$ rounds of iterative development. To further mitigate stochastic effects in competition outcomes, every entry in the scoring matrix $W$ is estimated by repeated matches. Detailed tournament prompts are listed in Appendix K.
We report the detailed scoring policy, generation configs, and repetition experiments in Appendix D. It is noteworthy that agents tend to generate different codes in repeated experiments, but their rankings are relatively stable.
4.2 Main Results
(a) Trends of global performance scores $G_{i}^{n}$.
(b) Quantitative comparison of learning abilities.
Figure 2: Visualization of agents’ learning patterns and scores.. For clarity, we use family names to represent LLM models instead of their full names. The results for other games are in Appendix E.
Table 3: Main LeaderBoard of CATArena. We conduct two groups of tournaments between minimal agents (rank from 1 to 6) and commercial agents (rank from 1 to 5), and report the average ranking across all four tasks. Metrics include S.C. (Strategy Coding), G.L. (Global Learning), and G.A. (Generalizability).
Agents Standard Variant G.A. $\downarrow$
S.C. $\downarrow$ G.L. $\downarrow$ S.C. $\downarrow$ G.L. $\downarrow$
Minimal
Claude-4-Sonnet 1.25 2.5 1.75 2.75 5.00
DeepSeek-Chat 5.75 2.75 4.25 2.75 2.75
Doubao-Seed 3.75 4.75 3.75 4.50 2.75
Gemini-2.5-Pro 3.25 3.75 3.25 2.75 3.25
GPT-5 3.75 3.50 3.00 3.75 2.25
Qwen3-Coder 2.25 3.75 3.00 4.5 4.75
Commercial
best ADK 3.25 2.25 2.00 3.75 2.50
Claude-Code 2.50 3.75 2.50 2.75 3.25
CodeX 2.25 2.75 3.00 3.00 3.25
Gemini-CLI 3.50 2.25 3.00 4.00 2.00
Qwen-Coder 3.00 3.75 4.00 1.25 3.25
Learning Ability.
Figure 2(a) visualizes the global performance scores $G_{i}^{n}$ revealing overall trends in agent strategies across iterations. Some agents, such as minimal agents driven by claude, exhibit a clear upward trajectory over multiple rounds, demonstrating strong learning capability. However, the performance of most agents remains unstable, and no obvious trend is observed. To explicitly illustrate the learning ability of each agent, we design a quantitative scoring method for global learning and introduce two additional learning modes: counter-adaption and self-improvement. The quantitative results for these three abilities are presented in Figure 2(b). Compared to Figure 2(a), these scores offer a more intuitive comparison of the agents’ learning strengths. This decomposition provides deeper insights into the mechanisms underlying agent learning. When both counter-adaption and self-improvement scores are positive, it indicates that the agent can effectively learn from both its opponents and itself, resulting in a positive global learning score. Compared to minimal agents, commercial agents exhibit stronger learning capabilities. Learning ability results for other games and a case study of strategy iteration mechanisms are provided in Appendix E. Through an analysis of the consistency of agent actions in endgame states, we observe that a majority of agents indeed learn from the code generated in the previous round, leading to increasing consistency. This phenomenon is most prominent in the Hold’em environment, possibly because its strategies are relatively simple and easier to learn.
Ability Evaluation.
We conduct two sets of tournaments: (1) minimal agents equipped with different LLMs, (2) the best-performing minimal agent against commercial code agents. For each tournament, we report the average ranking and scores for three core agent capabilities. The main leaderboard and main results of CATArena are summarized in Table 3 and Table 4. These results reveal the following key observations:
Observation 1: The performance gap among LLMs is more pronounced in minimal agents compared to commercial agents. Table 3 shows that Claude-4-Sonnet achieves the highest score among minimal agents, while the rankings of other LLMs are more dispersed. In contrast, commercial agents driven by the same LLMs exhibit much closer average rankings, with all agents scoring around 2.5 out of 5, indicating a reduced performance gap. Moreover, commercial agents demonstrate performance levels similar to the best-performing minimal agent. This suggests that the underlying agent framework can significantly influence how effectively an LLM’s capabilities are utilized, as commercial agents are often optimized for specific models.
Observation 2: The participating agents display different ranking orders across various capabilities. The tournament results reveal that the relative rankings of agents change depending on the specific core ability being tested. The ranking of these abilities provides a decomposition of end-to-end performance, offering insights for further optimization of both LLMs and agent frameworks.
Observation 3: Agents exhibit varied performance distributions across different tasks. The results indicate that agents’ performance is not uniformly distributed across all tasks, which is mainly attributed to the distinct nature and difficulty of the four tasks. In the variant tasks, the performance gap among agents is more pronounced, likely because the game rules and strategies are less familiar to agents.
[…]
Games Gomoku Hold’em Bridge Chess
Standard Variant Standard Variant Standard Variant Standard Variant
[…]
Minimal
Claude-4-Sonnet 0.88 -0.447 0.78 -0.156 -0.14 0.58 0.118 0.13 0.110 -0.45 0.79 0.174 1.0 0.047 0.005 0.90 -0.170 0.65 0.018 -0.55
Deepseek-Chat 0.23 0.027 0.38 0.077 0.13 0.01 0.010 0.00 -0.022 -0.01 0 0 0.10 0 0.10 0 0 0.10 0.049 0.10
Doubao-Seed 0.33 -0.192 0.72 -0.302 0.46 0.04 -0.035 0 0 -0.04 0.2 -0.033 0.45 -0.516 0.40 0.58 -0.337 0.10 0.034 -0.46
Gemini-2.5-Pro 0.25 -0.066 0.00 0.173 -0.12 0.01 0.020 0.00 0.078 -0.01 0.90 -0.195 0.60 -0.049 -0.30 0.58 -0.147 0.90 0.003 0.46
GPT-5 0.48 0.062 0.76 -0.019 0.18 0.16 0.102 0.87 -0.050 0.71 0.47 -0.095 0.10 0.293 -0.02 0.38 -0.525 0.45 -0 0.24
Qwen3-Coder 0.85 -0.523 0.36 -0.089 -0.50 0.20 0.038 0.00 0.003 -0.20 0.65 0.032 0.76 -0.230 -0.19 0.58 -0.187 0.80 -0.532 0.21
Commercial
best ADK 0 0.075 0.75 -0.022 0.88 0.07 0.073 0.46 0.030 0.39 0.25 0.295 0 0.361 -0.25 0.91 -0.110 1.00 -0.342 -0.47
Claude-Code 0.78 -0.322 0.66 0.194 -0.28 0.01 0.100 0 0.105 -0.001 1.00 -0.240 0.93 -0.139 -0.04 0.56 -0.158 0.44 -0.226 0.03
CodeX 0.47 0.454 0.69 -0.095 0.34 0.72 0.050 0.17 0.067 -0.55 0.75 0.098 0.50 0.285 -0.25 0.38 0.033 0.34 0.064 0.09
Gemini-CLI 0.31 0.260 0.19 0.172 0.0 0.13 0.050 0.37 -0.007 0.24 0.01 0.254 0.83 -0.286 0.79 0.38 0.395 0.38 -0.154 0.16
Qwen-Coder 0.94 -0.054 0.22 -0.530 -0.94 0.07 0.058 0 0.105 -0.007 0.49 0.157 0.25 0.556 -0.25 0.28 -0.204 0.34 0.039 0.19
4.3 Effectiveness of CATArena
We design a series of experiments to demonstrate the effectiveness of CATArena.
Comparison between Agents and LLM-Players.
The primary task in CATArena is strategy coding, which relies on the underlying coding capabilities of LLMs. We posit that reasoning over code to develop strategies is fundamentally different from direct reasoning during gameplay. To validate this distinction, we compare agent-developed strategies with the LLM-Player baseline (see Appendix F). Our results show that current agents primarily implement simple rule-based algorithms, indicating substantial room for advancement in agents’ strategy coding abilities. CATArena fully leverages this non-saturation, enabling sustainable iterative peer-learning.
To further analyze the similarities and differences between agent-implemented code strategies and those of LLM-Players, we ask agents’ code and LLM-Player to select the next action on endgame states. Figure 3 illustrates the action consistency between agents’ code and LLM-Players in Chess. Surprisingly, the strategies encoded in agent code differ significantly from those inferred directly by the LLM, even if they are from the same model. Meanwhile, strategies produced by different agents and different LLMs also show notable similarities. This indicates that strategy coding and reasoning in LLMs are distinct capabilities. We report results of other tasks are demonstrated in Appendix G. CATArena evaluates the strategy coding ability of agents rather than their reasoning ability, thereby filling a gap in previous benchmarks. The relationship between strategy coding ability and LLM-based strategy reasoning remains unclear and requires further investigation.
Figure 3: Action consistency between agents’ code and LLM-Players on Chess endgames.
Collective Learning Trends Among Agents.
Table 5: Collective learning trends of agents across different tasks in CATArena. $\text{DIS}_{\text{range}}$ and $\text{DIS}_{\text{std}}$ represent the Pearson correlation coefficients of the range and standard deviation of agent performance scores over four rounds, reflecting the similarity and dispersion of agent strategies. $\text{Trend}_{\text{mean}}$ denotes the Pearson correlation between the mean agent performance and the round number, indicating the overall trend of group improvement.
Gomoku Hold’em Bridge Chess
StdV. VarV. StdV. VarV. StdV. VarV. StdV. VarV.
$\text{DIS}_{\text{std}}$ -0.05 0.15 -0.81 -0.80 -0.82 -0.57 0.55 -0.04
$\text{DIS}_{\text{range}}$ -0.16 0.44 -0.80 -0.76 -0.54 -0.33 -0.08 0.16
$\text{Trend}_{\text{mean}}$ 0.42 -0.02 0.75 0.67 0.24 -0.10 -0.74 -0.79
We analyze the collective learning dynamics of agents across tasks, as presented in Table 5. The metrics $\text{DIS}_{\text{range}}$ and $\text{DIS}_{\text{std}}$ report the Pearson correlation between the standard deviation and range of agent performance scores over four rounds ($B_{i}^{n},n=1,2,3,4$). The higher similarity in performance scores (i.e., $\text{DIS}>0$) suggests that agents can learn effective strategies more readily, indicating lower task difficulty. Based on these results, the relative difficulty ranking of tasks in CATArena is Chess >> Gomoku >> Bridge >> Hold’em.
[…]
Additional Results.
CATArena’s iterative peer-learning framework is easily extensible to new tasks for evaluating other fundamental agent abilities. We demonstrate this by introducing a Machine Learning (ML) track and multi-lingual track, with experimental results provided in Appendix H and Appendix I, respectively. Experimental results indicate that current agents still exhibit substantial potential for improvement. As agents continue to advance, the open-ended task design and peer-learning evaluation framework of CATArena ensure that systematic assessment can be sustained over time. We also report agent cost in terms of token usage, time consumption, and generated code statistics for each agent, in Appendix J. Notably, Claude-4-Sonnet utilizes the most tools and tokens, and also develops a significantly larger amount of code. In contrast, GPT-5 achieves the best balance between token usage and performance. Efficient utilization of tokens and external tools remains an important research direction to advance the capabilities of LLM agents.

### v2

@@ 4 Experiments
In this section, we conduct comprehensive experiments to validate the effectiveness of the CATArena framework and analyze the evolutionary behaviors of code agents under the proposed iterative paradigm. Specifically, our evaluation aims to address the following four research questions:
  • RQ1: Can CATArena separate evolutionary capability ($S_{evo}$) from static base performance ($S_{base}$)?
  • RQ2: What dynamics do agents exhibit over iterations?
  • RQ3: How do peer-learning and self-reflection affect evolution?
  • RQ4: Is CATArena robust to task variants and extensible to new tracks?
Table 2: Main Results on CATArena. We report the Base Performance ($S_{\text{base}}$) and Evolutionary Capability ($S_{\text{evo}}$) of Minimal and Commercial Agents across six tasks. Best results in each column are bolded. Note: Commercial agents are excluded from SWE-Perf due to incompatibility with the Docker environment.
Agents Competitive Objective
Gomoku Hold’em Bridge Chess OIBench SWE-Perf
$S_{\text{base}}$ $S_{\text{evo}}$ $S_{\text{base}}$ $S_{\text{evo}}$ $S_{\text{base}}$ $S_{\text{evo}}$ $S_{\text{base}}$ $S_{\text{evo}}$ $S_{\text{base}}$ $S_{\text{evo}}$ $S_{\text{base}}$ $S_{\text{evo}}$
Minimal Claude 0.88 +0.030 0.58 +0.072 0.79 +0.062 0.90 -0.061 0.32 +0.040 0.02 -0.0016
DeepSeek 0.23 +0.014 0.01 +0.006 0.00 0.000 0.00 0.000 0.10 -0.029 0.02 -0.0011
Doubao 0.33 -0.046 0.04 -0.011 0.20 +0.013 0.58 -0.214 0.41 +0.041 0.00 +0.0002
Gemini 0.25 +0.156 0.01 +0.024 0.90 -0.083 0.58 -0.158 0.40 +0.022 0.01 +0.0012
GPT 0.48 +0.012 0.16 +0.025 0.47 -0.095 0.38 -0.044 0.41 +0.031 0.02 +0.0005
Qwen3 0.85 -0.196 0.20 +0.019 0.65 +0.003 0.58 -0.187 0.10 +0.099 0.01 +0.0036
Commercial Best ADK Agent 0.00 +0.176 0.07 +0.057 0.25 +0.191 0.91 -0.072 0.45 +0.049 - -
Claude Code 0.78 -0.035 0.01 +0.027 1.00 +0.056 0.56 -0.178 0.43 +0.072 - -
Codex 0.47 +0.103 0.72 -0.008 0.75 +0.062 0.38 +0.011 0.46 +0.186 - -
Gemini CLI 0.31 -0.030 0.13 -0.017 0.01 -0.056 0.38 +0.099 0.63 -0.026 - -
Qwen Code 0.94 +0.035 0.07 +0.025 0.49 +0.053 0.28 -0.144 0.35 +0.009 - -
@@ 4.1 Experimental Setups
We evaluate two categories of agents to disentangle model reasoning from system engineering: (1) Minimal Agents: To isolate the intrinsic evolutionary potential of foundation models, we implement a standardized minimal agent framework using Google Agent Development Kit (ADK) [^2] https://github.com/google/adk-python. This minimalist setup provides only core tools (e.g., file I/O and scripting), ensuring that performance differences are primarily attributable to the underlying LLM rather than complex scaffolding (Kim et al., 2023; Shihipar, 2026). We select a diverse set of API-based and open-weight models as backbones. (2) Commercial Code Agents: We include leading CLI-based commercial agents to establish high-performance industry baselines. These systems feature proprietary toolchains and represent the current state-of-the-art in agentic development. For brevity, we refer to agents by their model family names; full specifications are provided in Appendix D.
Evaluation Protocol.
All experiments are conducted within two distinct cohorts to facilitate intra-group evaluation: (1) A comparison among minimal agents equipped with various LLMs, and (2) A comparison between the best-performing minimal agent and a suite of commercial code agents. This setup ensures that agents are evaluated against peers within their respective categories to maintain a fair and reliable assessment of evolutionary capability.
Repetition and Stability. Due to the high computational and token costs of long-term evolution, the main evolutionary experiments are conducted as a single chain. To validate the reliability of this approach, we conducted a separate stability analysis on the initial rounds (Rounds 1 and 2), repeating the code generation process four times. We observe that while the specific code implementations vary across runs, the relative performance rankings among agents remain highly consistent (see Appendix E for analysis).
4.2 Results of Evolutionary Evaluation (RQ1)
Figure 2: Evolutionary Trajectories across Iterations. Solid lines represent the actual $G$ Score, dashed lines indicate the linear regression fits used to calculate $S_{\text{evo}}$.
Overall Performance Comparison. Table 2 summarizes the Base Performance ($S_{\text{base}}$) and Evolutionary Capability ($S_{\text{evo}}$) across all agents. We also report statistics on HTTP error rates in Appendix F. The results yield three key observations: (1) We observe a weak correlation between $S_{\text{base}}$ and $S_{\text{evo}}$ across all tasks. High $S_{\text{base}}$ does not guarantee effective evolution. For instance, while Claude-4 dominates the initial round in Chess ($0.90$), it suffers significant regression (0.061-0.061) during evolution. Conversely, Gemini-2.5-Pro demonstrates a low-start, high-growth trajectory in Gomoku ($0.25\rightarrow+0.156$). While it starts with a deficit, its high $S_{\text{evo}}$ indicates the ability to rapidly absorb feedback. (2) Evolutionary dynamics are inherently task-dependent. We observe a distinct complexity barrier in the evolution of agents. In high-difficulty environments such as Chess and SWE-Perf, most agents exhibit negligible or negative $S_{\text{evo}}$. This indicates that when the reasoning gap between the current implementation and the optimal solution exceeds a certain threshold, agents struggle to extract actionable insights from feedback. (3) Commercial frameworks primarily enhance static generation rather than evolution ability. While commercial agents frequently outperform the Best ADK Agent in $S_{\text{base}}$, they do not exhibit a decisive advantage in $S_{\text{evo}}$. This disparity underscores that evolutionary capability is a distinct dimension of intelligence, largely unaddressed by current commercial scaffolding, and represents a critical frontier for future agentic system research.
Evolutionary Trajectories. Figure 2 visualizes the round-wise performance dynamics of global performance $G^{n}$, highlighting three critical phenomena. First, we observe significant performance oscillation across iterations. This volatility indicates that current agents lack robust self-evaluation mechanisms, often discarding valid strategies for inferior ones during the evolutionary process. Second, despite this instability, evolution can drive substantial performance leaps. For instance, in OIBench, Qwen (blue line) advances from a last-place initialization ($G_{1}=0.10$) to the top-performing tier ($G_{4}=0.43$) through iterative refinement. This trajectory confirms that static zero-shot benchmarks may underestimate an agent’s true potential, which can be significantly amplified through iterative feedback. Third, the visualization validates the reliability of our metric design. As indicated by the dashed regression lines, $S_{\text{evo}}$ effectively captures the global optimization trend, proving robust to transient performance drops caused by exploratory failures typical of stochastic code generation. The results for other tasks are detailed in Appendix G.
4.3 Temporal Dynamics of Evolution (RQ2)
We investigate the temporal dynamics of agent evolution by analyzing population statistics across task types and extending the evolutionary horizon to $N=7$ rounds.
Population Evolution Patterns. Table 3 characterizes the evolutionary trends using two metrics derived from Pearson correlation coefficients: $\text{Trend}_{\text{mean}}$, representing the correlation between round number and average population performance, and $\text{DIS}_{\text{std}}$, the correlation between round number and score dispersion (standard deviation). The results reveal that the population structure is heavily dictated by the feedback mechanism. Objective optimization tasks tend to drive divergence ($\text{DIS}_{\text{std}}>0$), whereas competitive tasks foster convergence ($\text{DIS}_{\text{std}}<0$). This disparity arises because objective feedback rewards strong agents capable of approaching theoretical limits, amplifying the gap, while the open-source nature of competitive tournaments allows weaker agents to mimic successful strategies, homogenizing the population. Regarding overall performance, the majority of tasks exhibit a positive $\text{Trend}_{\text{mean}}$, confirming that iterative refinement generally drives collective improvement. The only exception is Chess ($\text{Trend}_{\textbf{mean}}<0$), where the excessive reasoning complexity overwhelms the feedback loop, leading to a collapse rather than evolution.
Table 3: Evolutionary Population Dynamics. Pearson correlation coefficients between iteration rounds and population performance metrics: $\text{DIS}_{\text{std}}$ (Standard Deviation) and $\text{Trend}_{\text{mean}}$ (Mean Score).
Task Type Task $\text{DIS}_{\text{std}}$ $\text{Trend}_{\text{mean}}$ Pattern
Objective OIBench +0.42 +0.79 Divergence
SWE-Perf +0.20 +0.20 Weak trend
Competitive Gomoku -0.05 +0.42 Convergence
Hold’em -0.81 +0.75 Strong convergence
Bridge -0.82 +0.24 Moderate convergence
Chess +0.55 -0.74 Oscillation
Figure 3: Long-term Evolutionary Dynamics ($N=7$). Top: The Peak Performance Gain $\Delta\text{Max}$ across iterations. Bottom: The trajectories of Global Performance $G^{n}$ for individual agents.
Long-term Evolution. We extend the evaluation to $N=7$ for Gomoku and OIBench to investigate long-term evolution, visualizing the results in Figure 3. We first analyze the Peak Performance Gain ($\Delta\text{MAX}$), defined as the sum of performance increments across all models between consecutive rounds, which quantifies the evolutionary vitality. We observe that $\Delta\text{MAX}$ is significantly higher during the initial rounds ($N<4$) compared to the later stages ($N>4$). This decay justifies our selection of a four-round window for the main experiment, as it effectively captures the most active phase of evolution. Furthermore, the global performance $G^{n}$ trajectories in the bottom panel reveal a clear pattern of stratification. While early rounds exhibit frequent rank crossings, the rankings stabilize in the long term with significantly fewer intersections. Most agents eventually trend towards saturation, failing to break through performance ceilings in the later rounds. This suggests that while iterative feedback optimizes strategies, the asymptotic performance remains bounded by the intrinsic reasoning limits of the underlying models.
4.4 Mechanism of Evolution (RQ3)
To deconstruct the black box of agent evolution, we categorize the evolutionary process into two distinct mechanisms: Self-Reflection, where the agent introspectively analyzes its own execution logs to fix bugs and refine logic; and Peer-Learning, where the agent extrospectively analyzes external solutions to adopt superior strategies or algorithms.
Table 4: Ablation Study on peer-learning. Comparison of performance with full feedback ($G_{2}$) versus self-reflection only ($G_{2}^{\prime}$). Bold values indicate improvement over the initial round ($G_{1}$). The Pattern column denotes the effective mechanism: S indicates effective Self-Reflection ($G_{2}^{\prime}>G_{1}$), P indicates effective Peer-Learning ($G_{2}>G_{2}^{\prime}$).
Agent Gomoku OIBench
$G_{1}$ $G_{2}^{\prime}$ $G_{2}$ Pattern $G_{1}$ $G_{2}^{\prime}$ $G_{2}$ Pattern
Claude 0.814 0.520 0.951 ✗P 0.318 0.204 0.410 ✗P
DeepSeek 0.363 0.265 0.333 ✗✗ 0.097 0.105 0.000 S✗
Doubao 0.716 0.039 0.598 ✗✗ 0.408 0.458 0.000 S✗
Gemini 0.225 0.794 0.118 S✗ 0.403 0.346 0.301 ✗✗
GPT 0.500 0.480 0.020 ✗✗ 0.413 0.470 0.423 S✗
Qwen 0.833 0.814 0.618 ✗✗ 0.101 0.084 0.398 ✗P
Ablation Study. To isolate the contribution of two evolution mechanisms, we compare agent performance with full feedback ($G_{2}$) against a self-reflection-only baseline ($G_{2}^{\prime}$). We omit the symmetric ablation (peer-only without logs) as it reduces agents to blind imitation lacking performance grounding. As detailed in Table 4, the results reveal a critical limitation: current agents struggle to synergize peer code and execution logs. Instead of benefiting from a unified feedback loop, agents typically exhibit a dependency on a single source. Specifically, the majority of models still rely predominantly on self-reflection for improvement, often treating peer information as noise or distraction. Only select high-capability agents (Claude and Qwen) demonstrate the ability to effectively leverage peer-learning to achieve performance breakthroughs. This lack of synthesis capability points to a key direction for future research: developing mechanisms that enable agents to robustly reconcile internal diagnostics with external knowledge.
Code Consistency Analysis. To further dissect the behavioral mechanisms underlying evolution, we quantify the extent to which agents rely on self-refinement versus peer-adoption. For each agent’s Round 2 submission, we compute its similarity to its own Round 1 code (X-axis, self-consistency) and the maximum similarity to any peer’s Round 1 code (Y-axis, peer-adoption). For competitive tasks, similarity is measured by behavioral alignment in endgame positions; for objective tasks, we employ CodeBLEU (Ren et al., 2020) to assess syntactic and semantic overlap. As illustrated in Figure 4, most points cluster near the diagonal, indicating a balanced use of both mechanisms. However, three extreme failure modes emerge: (1) Blind Copying (high $y$), where Doubao and DeepSeek in Hold’em completely abandon their logic to mimic peers; (2) Stubborn Stagnation (high $x$), exemplified by DeepSeek in Gomoku, which refuses to adapt despite poor performance; and (3) Chaotic Reconstruction ($x,y<20$), where agents discard both internal and external priors, leading to unstable solutions. These extremes highlight a critical inability of current models to dynamically balance exploration and exploitation. To provide a more granular view of these behavioral patterns, we present the complete pairwise similarity matrices for all tasks in Appendix H.
Figure 4: Code Evolution Patterns. Each point represents an agent’s Round 2 submission. The X-axis measures similarity to its own Round 1 code, while the Y-axis measures the maximum similarity to any peer’s Round 1 code.
Case Study. Development logs reveal a structured refinement process driven by peer-learning and self-reflection (See Appendix I). Agents actively utilize tools such as Grep to extract win statistics from tournament reports, identifying top-performing peers to study and synthesize their logic. For instance, in Texas Hold’em, Claude transitioned from a basic heuristic formula to a sophisticated opponent-aware strategy after analyzing the aggressive betting patterns of high-ranking competitors. This peer-learning is complemented by explicit self-reflection, where agents benchmark their own win rates against rivals and plan targeted improvements in modules such as risk management or search depth. These behaviors show that agents evolve through deliberate optimization rather than stochastic mutations.
4.5 Robustness and Extensibility of CATArena (RQ4)
Generalization to Unseen Variants. To verify that CATArena evaluates genuine adaptive capabilities rather than the retrieval of memorized solutions, we conducted experiments on unseen variants for five tasks. These variants introduce rule perturbations, such as inverting hand rankings in Hold’em, that invalidate standard strategies present in the agents’ pre-training data. Detailed results are provided in Appendix J. As shown in Table 11, while the variance in zero-shot base scores $S_{\text{base}}$ increases significantly due to non-robust agents failing to generate valid strategies, the evolutionary scores $S_{\text{evo}}$ maintain their discriminative power. This consistency confirms that CATArena robustly measures the intrinsic evolutionary ability of agents, independent of data contamination or specific task familiarity.
Extensibility to Diverse Tracks. CATArena is designed as a modular framework, allowing seamless integration of new environments and evaluation tracks. To demonstrate this extensibility, we incorporated three additional experimental tracks: a Multilingual Track (Python, C++, Java, JS, Go), a Machine Learning (ML) Track, and a Complex Task Track (Pommerman). Detailed results are provided in Appendices K, L, and M, respectively. We observe that advanced tracks pose significant challenges for SoTA agents. Most agents exhibit low $S_{\text{base}}$ and unstable evolutionary trajectories, similar to the oscillation patterns observed in Chess. This suggests that while the framework is ready for broader evaluation, current models struggle to generate robust starting points for evolution in these domains. As agent capabilities advance, CATArena will scale to these harder tasks to continue evaluating the frontier agents.